OpenStudy (jinx.exe):

What is the interquartile range of the data? 150, 165, 180, 198, 204, 216, 230, 255 A. 105 B. 51 C. 50.5 D. 28.5

OpenStudy (igreen):

IQR = Q3 - Q1

OpenStudy (anonymous):

180, 198, 204, 216

OpenStudy (igreen):

150, 165, 180, 198, 204, 216, 230, 255 Cut the data in half 150, 165, 180, 198 and 204, 216, 230, 255 Find the middle terms. 150, 165, 180, 198 The two middle terms are 165 and 180, so find the mean. 165 + 180 = ?

OpenStudy (anonymous):

is my answer but igreen might have a different answer

OpenStudy (igreen):

Can you add 165 + 180? @Jinx.exe

OpenStudy (jinx.exe):

is the answer 105..? That's what I've got.

OpenStudy (jinx.exe):

165+180=345

OpenStudy (igreen):

No, 105 is incorrect. Now divide 345 / 2

OpenStudy (jinx.exe):

172.5?

OpenStudy (igreen):

Yes, so that's Q1, now let's find Q3. 204, 216, 230, 255 The two numbers in the middle are 216 and 230. Add 216 + 230 then divide 2.

OpenStudy (jinx.exe):

223..?

OpenStudy (igreen):

Correct, so Q3 is 223. Now subtract 223 - 172.5

OpenStudy (jinx.exe):

50.5?

OpenStudy (igreen):

Correct, that's the Interquartile Range.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'. @Jinx.exe